OpenStudy (anonymous):

Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days.

OpenStudy (anonymous):

I was just tagging you! Haha

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Okay... let me think

OpenStudy (anonymous):

Ah, think I got it

OpenStudy (anonymous):

Okay. Explain how you got it? :)

OpenStudy (anonymous):

Oh, wait, misread the question

OpenStudy (anonymous):

Do you know what standard form is?

OpenStudy (anonymous):

f(x) = ... right?

OpenStudy (anonymous):

It's like, a + b + c + d = 0, everything on one side and then set it to 0

OpenStudy (anonymous):

Ah. Okay.

OpenStudy (anonymous):

So f(5) = 465 is like f(5) = 93(5) which is like f(x) = 93x

OpenStudy (anonymous):

I think. o.o

OpenStudy (anonymous):

Lets solve for x. f(5) = 465 f(7) = 625

OpenStudy (anonymous):

We have to reverse this, I think. something like that. I'm confuzzled

OpenStudy (anonymous):

Ditto broski. xD

OpenStudy (anonymous):

I think that this problem has two solutions.

OpenStudy (anonymous):

Is this problem from flvs?

OpenStudy (anonymous):

Maybe. 0_0

OpenStudy (anonymous):

No it's fine, this is tutoring, I need to know what lesson. Algebra II?

OpenStudy (anonymous):

.....algebra1...

OpenStudy (anonymous):

Oh yeah, functions are in algebra I... hmmz

OpenStudy (anonymous):

okay, I made a chart on paper, when x = 5, y = 465 when x = 7, y = 625

OpenStudy (anonymous):

lets find the slope so I can graph it out

OpenStudy (anonymous):

Okay. \[\frac{ 5 - 7 }{ 465 - 625 }?\]

OpenStudy (anonymous):

465 - 625 over 5 - 7 -160 over -2 80

OpenStudy (anonymous):

it's y2 - y1 over x2 - x1

OpenStudy (anonymous):

Now I can fill in my other points and graph it, gimme a sec

OpenStudy (anonymous):

Ohh. Okay

OpenStudy (anonymous):

standard form for a linear equation is Ax + By = C

OpenStudy (anonymous):

Uhm. okay. 0_0

OpenStudy (anonymous):

>.< so hard this one is

OpenStudy (anonymous):

Right. xD

OpenStudy (anonymous):

I can come back to this one later. Haha. Wanna move on?

OpenStudy (amistre64):

Ax + By = C converts to: y = -A/B x + C/B

OpenStudy (anonymous):

no we can do this

OpenStudy (anonymous):

I feel like the answer is staring at me.

OpenStudy (anonymous):

65 = y intercept...

OpenStudy (amistre64):

since you know the slope, and a given point: use the point slope form: y-y1 = m(x-x1) convert it to slope intercepts as y = mx +(-mx1) + y1 C/B = -A/B x1 + y1, solving for C we get C = -Ax1 + By1 which is practically the end results, but they dont like a -A

OpenStudy (anonymous):

y = mx + b y = (80)x + 65 y = 80x + 65 standard form is... -80x + y = 65

OpenStudy (anonymous):

but we don't want a -a

OpenStudy (amistre64):

almost ... but needs a slight modification

OpenStudy (anonymous):

Uhhhh?

OpenStudy (anonymous):

so we multiply everything by -1.

OpenStudy (amistre64):

good job :)

OpenStudy (anonymous):

80x - y = -65! :D

OpenStudy (anonymous):

wow I feel good. tyvm amistre64

OpenStudy (anonymous):

I'm kinda lost.

OpenStudy (anonymous):

Okay let me type up an explanation

OpenStudy (anonymous):

The slope-intercept form of a line is: y = mx + b m = slope and b = y intercept y = 80x + 65 Standard form is Ax + By = C A must be positive, A and B cannot both be 0, and A, B, And C must be integers. y = 80x + 65 -80x + y = 65 Now we have to multiply it all by -1 to make a positive -1(-80x + y = 65) 80x - y = -65

OpenStudy (anonymous):

Do you get it now?

OpenStudy (anonymous):

@RockerSk8er

OpenStudy (anonymous):

I'm here? Yes I get it! thank you so much! :D

OpenStudy (anonymous):

No problem buddy. :) Got any more?

OpenStudy (anonymous):

Yeah, lemme put the answer in first. :) These are essay questions so. Ugh.

OpenStudy (anonymous):

Mk

OpenStudy (anonymous):

What equation would be my final answer?

OpenStudy (anonymous):

80x - y = -65

OpenStudy (anonymous):

x is the number of days, so if he rented it for two days, it would be 160 - y = -65 -y = 95 y = -95 dollars

OpenStudy (anonymous):

wait lol oops

OpenStudy (anonymous):

it would be y = -225

OpenStudy (anonymous):

gotta subtract it, not add it. but yeah, final equation is 80x - y = -65

OpenStudy (anonymous):

grr, y = 225 I stupid.

OpenStudy (anonymous):

I was looking at it, and didn't get why it equals -65?

OpenStudy (anonymous):

lerl I messed up, it's 160 - y = -65 160 + (-160) - y = -65 + (-65) -y = -225 y = 225

OpenStudy (anonymous):

Okay, I see now.

OpenStudy (anonymous):

but 80x - y = -65 is in standard form, which you want post more questions if you want more halp. :D

OpenStudy (anonymous):

"Write the equation obtained in Part A using function notation."

OpenStudy (anonymous):

post as a new question so we stop spamming amis

OpenStudy (anonymous):

Ah, okay.